I'm having a hard time with this problem. Please help!
2) For the function y = x2 - 6x + 8, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer:

Found 2 solutions by Nate, stanbon:
Answer by Nate(3500)

y = x^2 - 6x + 8
y - 8 = x^2 - 6x
y - 8 + 9 = (x - 3)^2
y + 1 = (x - 3)^2
y = (x - 3)^2 - 1

Answer by stanbon(75887)

For the function y = x^2 - 6x + 8, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
--------------
Rewrite as:
y-8 = x^2-6x
Complete the square, as follows:
x^2-6x+9=y-8+9
(x-3)^2=y+1
y=(x-3)^2-1
Cheers,
Stan H.